OpenStudy (anonymous):

a-b=p and a =b=k then a^2-b^2=?

OpenStudy (anonymous):

u mean if a-b=p and a+b=k then a^2-b^2?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok this is a little hint for u : difference of 2 squares :\[a^2-b^2=(a-b)(a+b)\]